# Proactive and Reactive Constraint Programming for Stochastic Project Scheduling with Maximal Time-Lags

### v1

@@ Abstract
This study investigates scheduling strategies for the stochastic resource-constrained project scheduling problem with maximal time lags (SRCPSP/max)). Recent advances in Constraint Programming (CP) and Temporal Networks have re-invoked interest in evaluating the advantages and drawbacks of various proactive and reactive scheduling methods. First, we present a new, CP-based fully proactive method. Second, we show how a reactive approach can be constructed using an online rescheduling procedure. A third contribution is based on partial order schedules and uses Simple Temporal Networks with Uncertainty (STNUs). Our statistical analysis shows that the STNU-based algorithm performs best in terms of solution quality, while also showing good relative offline and online computation time.
@@ 2 The Scheduling Problem
The RCPSP/max problem [KS96] is defined as a set of activities $J$ and a set of resources $R$, where each activity $j\in J$ has a duration $d_{j}$ and requires from each resource $r\in R$ a certain amount indicated by $r_{r,j}$. A solution to the RCPSP/max is a start-time assignment $s_{j}$ for each activity, such that the following constraints are satisfied: (i) at any time the number of resources used cannot exceed the max resource capacity $c_{r}$; and (ii) for some pairs $(i,j)$ of activities, precedence constraints are defined as minimal or maximal time lags between the start time of $i$ and the start time of $j$. The goal is to minimize the makespan of the schedule.
We use the following example from [SFSW13] of a simple RCPSP/max instance. There are five activities: $a$, $b$, $c$, $d$, and $e$, with start times $s_{a}$, $s_{b}$, $s_{c}$, $s_{d}$, and $s_{e}$. The activities have durations of 2, 5, 3, 1, and 2 time units, respectively, and resource requirements of 3, 2, 1, 2, and 2 on a single resource with a capacity of $c_{r}=4$. The precedence relations are as follows: activity $b$ starts at least 2 time units after $a$ starts, activity $b$ starts at least 1 time unit before $c$ starts, activity $c$ cannot start later than 6 time units after $a$ starts, and activity $d$ starts exactly 3 time units before $e$ starts. These constraints can be visualized in a project graph (Figure 1). A feasible solution to this problem is $\{s_{a}=1,s_{b}=3,s_{c}=4,s_{d}=0,s_{e}=3\}$, for which the Gannt chart is visualized in Figure 2 (ignore the arrows for now).
Figure 1: Example of project graph, figure from [SFSW13].
A special property of (S)RCPSP/max is the following:
@@
This property can be helpful when reusing start time solutions for a mutated instance, for example, because of stochastic activity durations. SRCPSP/max is an extension in which the durations follow a stochastic distribution [FLVX12]. The durations $d_{j}$ are random variables bounded by an uncertainty interval and will become available when an activity finishes. Earlier work on this stochastic variant introduces the $\alpha$-robust makespan [FLVX12], that is the expected makespan value for a scheduling strategy for which the probability that the schedule is feasible is at least $1-\alpha$, as a metric for comparing different methods.
@@ 3.1.2 SORU and SORU-H
Since $\mathsf{SORU}$ is computationally expensive, the authors also propose a heuristic version, dubbed $\mathsf{SORU}$-$\mathsf{H}$. Instead of a set of samples, one single summarizing sample is used, which represents a quantile of the distribution. Note that because of Proposition 1, this heuristic approximates the $\alpha$-robust makespan. At the same time, it is much cheaper to compute because typically the runtime increases for a larger sample size in SAA, as shown by [VFL16].
Since nowadays CP is the state of the art for deterministic project scheduling [NRR23], we re-investigate SAA approaches for SRCPSP/max with CP. We expect that the scalability of SAA is much less of a problem than it was when the $\mathsf{SORU}$ methods were presented.
@@ 3.2.1 Constructing Ordering Constraints
A partial order schedule (POS) can be seen as a collection of schedules that ensure resource feasibility, but maintain temporal flexibility. A POS is defined as a graph where nodes represent activities, and edges temporal constraints between them. There are several methods to derive a POS. In the original paper by [PSCO04], two approaches are outlined to construct the ordering constraints between activities, either analyzing the resource profile to avoid all possible resource conflicts, based on Minimal Critical Sets (MCS) [LR83], or using a fixed resource-feasible schedule together with a chaining procedure to construct resource chains (see Figure 2). They find that using the single-point solution together with chaining seems both simple and most effective. Subsequent work explored alternative MCS-based approaches [LM09, LMB13] and chaining heuristics [FLVX12].
Figure 2: Example Gannt chart figure adjusted from [FLVX12]. The X-axis is time, and Y-axis shows resource demand. The arrows indicate an example of generating a POS with chaining, starting from a fixed solution schedule.
@@ 3.2.3 Execution Strategies
An STNU is dynamically controllable (DC) if there is a strategy to determine execution times for all controllable (non-contingent) time points which ensures that all temporal constraints are met, regardless of the outcomes of the contingent links. Some DC-checking algorithms generate new so-called wait edges to make the network DC [Mor14]. The DC STNU with wait edges is referred to as an Extended STNU (ESTNU) and can be given to a Real-Time Execution Algorithm. Such an algorithm is the online component that transforms an STNU into a schedule (i.e. an execution time for each time point), given the observations for the contingent time points. An algorithm specifically tailored to ESTNUs is RTE [HP24].
@@ 3.3 Benchmarking Approaches
Benchmarking procedures for comparing scheduling methods are inconsistent in the literature, with varying problem sets and comparison methods. Some studies used industrial scheduling instances [LM09, LMB13], while the majority [PSCO04, FLVX12, VFL16, FVL16] relied on PSPlib [KS96] instances, which are deterministic and transformed into stochastic versions with noise. Research typically focused on instances with 10, 20, and 30 activities (j10-j30). Different studies assessed varying metrics, such as schedule flexibility and robustness [PSCO04, PCOS07], solver performance [LM09, LMB13], or the $\alpha$-robust makespan [FLVX12, VFL16, FVL16]. However, no comprehensive benchmarking paper exists that evaluates both solution quality and computation time while also correctly accounting for infeasibilities. We take inspiration from [LF03], who compare different planners that can fail, providing a framework for comparing solution quality and speed while also correctly considering correctly considering these failures.
@@ 4 Scheduling Methods
This section outlines the proposed methods. We explain how to use CP for these scheduling problems so far dominated by MIP approaches.
@@ 4.1 Constraint Programming for RCPSP/max
$\sum_{j\in J}\mathrm{Pulse}(x_{j},r_{r,j})\leq c_{r}$ $\ \forall r\in R$ (1e)
$x_{j}:\mathrm{IntervalVar}(J,d_{j})$ $\ \forall j\in J$ (1f)
For the deterministic RCPSP/max, we use the modern interval constraints from the IBM CP optimizer [te17]. The CP model is defined in equations 1a-1f , we use IBM’s syntax and modify the RCPSP example from [LRSV18] to RCPSP/max. We use the earlier introduced nomenclature together with the minimal time lags $min_{j,i}$ and maximal time lags $max_{j,i}$ that are the temporal differences between start times of activities $j$ and $i$ if $i$ is a successor of $j$. We introduce the decision variable $x_{j}$ as the interval variable for activity $j\in J$. The $\mathrm{Pulse}$ function generates a cumulative expression over a given interval $x_{j}$ with a certain value. For a task $j$, this value is its resource usage $r_{r,j}$. The aggregated pulse values are constrained so that their sum does not exceed the total available resource capacity $c_{r}$.
@@ 4.2.1 Proactive Method
For this SAA method, we can reuse the deterministic RCPSP/max CP model, introduced in Section 4.1, but we introduce scenarios. This model is inspired by the MIP version by [VFL16]. A special variant is the SAA with only one sample, for which a $\gamma$-quantile can be used which we call $\mathsf{proactive}_{\gamma}$. If a feasible schedule can be found for the $\gamma$-quantile, this schedule will also be feasible for all duration realizations on the left-hand side of the $\gamma$-quantile because of Proposition 1. We provide the SAA model below. We use the same nomenclature as for the deterministic model, but we introduce the notion of scenarios $\omega\in\Omega$, and find a schedule $s_{j}\ \forall j\in J$ that is feasible for all scenarios it has seen if one exists.
@@
$\sum_{j\in J}\mathrm{Pulse(x_{j}^{\omega},r_{r,j})\leq c_{r}}$ $\ \forall r\in R\ \forall\omega\in\Omega$ (2e)
$x_{j}^{\omega}:\mathrm{IntervalVar(J,y_{j}^{\omega})}$ $\ \forall j\in J\ \forall\omega\in\Omega$ (2f)
$s_{j}==\mathrm{start}(x_{j}^{\omega})$ $\ \forall j\in J\ \forall\omega\in\Omega$ (2g)
@@ 4.2.2 Reactive Method
Fully reactive approaches, involving complete rescheduling by solving a deterministic RCPSP/max, have been considered impractical due to high computational demands and low schedule stability [VdVDH07]. However, advances in CP for scheduling [LRSV18, NRR23] mitigate these issues. In practice, decision-makers often reschedule their entire future plan when changes occur. Thus, including this reactive approach in our comparison is valuable. To our knowledge, such an approach has not been evaluated before. The outline is:
@@
  • At every decision moment (i.e. when an activity finishes) resolve the deterministic RCPSP/max while fixing all variables until the current time to reschedule with new information, we again use the estimation $\hat{d}$ for the activities that did not finish yet. Resolving is only needed when the finish time of an activity deviates from the estimated finish time. We warm start the solver with the previous solution.
@@ 4.2.3 STNU-based Method
The resulting STNU is tested for dynamic controllability (DC), and if the network is DC its extended form (ESTNU) is given to the RTE algorithm. Since makespan minimization is of interest, we slightly adjust the algorithm from [HP24]: instead of selecting an arbitrary executable time point and an arbitrary allowed execution time, we always choose the earliest possible time point at the earliest possible execution time.
@@ 4.3 Statistical Tests for Pairwise Comparison
A strategy for benchmarking is to provide partial orderings of the scheduling methods for the different metrics of interest (i.e. solution quality, runtime offline, runtime online). A partial ordering can be obtained by executing pairwise comparisons of the methods per problem size and per metric, taking inspiration from [LF03].
@@
An alternative test to the Wilcoxon test that is also used by [LF03] is the proportion test (see Test 4 in the book by [Kan06]). This test is weaker than the Wilcoxon, but provides at least information about significance in the proportion of wins when the Wilcoxon test shows no significant difference.
An interesting additional test is to test whether there is a significant difference in the magnitude of the different metrics. This can be tested with a pairwise t-test on two related samples of scores (see Test 10 in the book by [Kan06]). This test can only be performed on double hits, because infinitely bad computation time or solution quality will disturb the test.
@@
Table 1: Pairwise test results for the instance set ubo50, noise level $c=1$, covering solution quality, offline and online runtime. For each metric, we show the test results of the Wilcoxon test, proportion test, and magnitude test such as described in Section 4.3. Wilcoxon test results are presented as: [nr pairs] z-value (p-value), with (*) for $p<0.05$. Proportion test results are: [nr pairs] proportion (p-value). Magnitude test results are: [nr pairs] t-stat (p-value), with normalized averages of the two methods in the last two rows. Additional results for $c\in[1,2]$, j10, j20, j30, ubo50, and ub0100 are in the Technical Appendix. Each column header lists the two methods compared, with the first being the better performer, excluding pairs with only zero-differences.
@@ 5 Experimental Evaluation
The goal of the evaluation is to analyze the relative performance of the different scheduling methods proposed in our methodology.
@@ 5.1 Data Generation
We use the j10, j20, j30, ubo50, and ubo100 sets from the PSPlib [KS96]. Instead of using all instances from j10 to j30, we select 50 per set and extend previous work [VFL16, FVL16] by including 50 instances each from the ubo50 and ubo100 sets. Following prior research, we use deterministic durations to set up stochastic distributions with different noise levels, converting deterministic instances into stochastic ones. We use uniform discrete distributions for durations based on the deterministic processing times $d_{j}$. Specifically, we define the lower bound as $lb=\max(1,\text{round}(d_{j}-c\cdot\sqrt{d_{j}}))$ and the upper bound as $ub=\text{round}(d_{j}+c\cdot\sqrt{d_{j}})$, where we vary the noise level $c=\{1,2\}$. The source and sink nodes always have a deterministic duration of zero. Each evaluation of a method corresponds to one sample from the distribution, and we sample $10$ times for each instance. We excluded the instance samples for which it is impossible to find a feasible schedule (i.e. the deterministic problem with perfect information is infeasible for the given activity duration sample).

### v2

@@ Abstract
This study investigates scheduling strategies for the stochastic resource-constrained project scheduling problem with maximal time lags (SRCPSP/max)). Recent advances in Constraint Programming (CP) and Temporal Networks have re-invoked interest in evaluating the advantages and drawbacks of various proactive and reactive scheduling methods. First, we present a new, CP-based fully proactive method. Second, we show how a reactive approach can be constructed using an online rescheduling procedure. A third contribution is based on partial order schedules and uses Simple Temporal Networks with Uncertainty (STNUs). Our analysis shows that the STNU-based algorithm performs best in terms of solution quality, while also showing good relative computation time.
@@ 2 The Scheduling Problem
The RCPSP/max problem [KS96] is defined as a set of activities $J$ and a set of resources $R$, where each activity $j\in J$ has a duration $d_{j}$ and requires from each resource $r\in R$ a certain amount indicated by $r_{r,j}$. A solution to the RCPSP/max is a start-time assignment $s_{j}$ for each activity, such that the constraints are satisfied: (i) at any time the number of resources used cannot exceed the max resource capacity $c_{r}$; and (ii) for some pairs $(i,j)$ of activities, precedence constraints are defined as minimal or maximal time lags between the start time of $i$ and the start time of $j$. The goal is to minimize the makespan.
We use the following example from [SFSW13] of a simple RCPSP/max instance. There are five activities: $a$, $b$, $c$, $d$, and $e$, with start times $s_{a}$, $s_{b}$, $s_{c}$, $s_{d}$, and $s_{e}$. The activities have durations of 2, 5, 3, 1, and 2 time units, respectively, and resource requirements of 3, 2, 1, 2, and 2 on a single resource with a capacity of $c_{r}=4$. The precedence relations are as follows: activity $b$ starts at least 2 time units after $a$ starts, activity $b$ starts at least 1 time unit before $c$ starts, activity $c$ cannot start later than 6 time units after $a$ starts, and activity $d$ starts exactly 3 time units before $e$ starts. These constraints can be visualized in a project graph (Figure 1). A feasible solution to this problem is $\{s_{a}=1,s_{b}=3,s_{c}=4,s_{d}=0,s_{e}=3\}$, for which the Gannt chart is visualized in Figure 2.
abcde0000333333-3- 322222222111166-6- 62222
Figure 1: Example project graph [SFSW13].
A special property of RCPSP/max is the following:
@@
This property can be helpful when reusing start times for a mutated instance, for example, because of stochastic activity durations. The Stochastic RCPSP/max (SRCPSP/max) is an extension in which the durations follow a stochastic distribution [FLVX12]. We define the problem as follows: the activity durations are independent random variables indicated by $d_{j}$ representing the duration of activity j. We assume that each $d_{j}$ follows a discrete uniform distribution, $d_{j}\sim$ DiscreteUniform($lb_{j},ub_{j}$) where $lb_{j}$, $ub_{j}$ denote the minimum and maximum possible durations for activity $j$. Each duration becomes available when an activity finishes.
In general, comparing methods for this problem is not straightforward as while comparing method A and B, it could be that only method A obtains a feasible solution. Looking at double hits (instances solved by both methods) is a reasonable choice, but neglects that potentially one of the two methods solved more problem instances than the other. Earlier work on this stochastic variant introduces the $\alpha$-robust makespan [FLVX12], that is the expected makespan value for a scheduling strategy for which the probability that the schedule is feasible is at least $1-\alpha$, as a metric for comparing different methods but also indicated the limitations of this metric.
@@ 3.1.2 SORU and SORU-H
Since $\mathsf{SORU}$ is computationally expensive, the authors propose a heuristic version, dubbed $\mathsf{SORU}$-$\mathsf{H}$. Instead of a set of samples, one summarizing sample is used, which represents a quantile of the distribution. Note that because of Proposition 1, this heuristic approximates the $\alpha$-robust makespan. At the same time, it is much cheaper to compute because typically the runtime increases for a larger sample size in SAA, as shown by [VFL16].
Since nowadays CP is the state of the art for deterministic project scheduling [NRR23], we re-investigate SAA approaches for SRCPSP/max with CP and compare this CP-based proactive approach to reactive approaches.
@@ 3.2.1 Constructing Ordering Constraints
A partial order schedule (POS) can be seen as a collection of schedules that ensure resource feasibility, but maintain temporal flexibility. A POS is defined as a graph where nodes represent activities, and edges temporal constraints between them. There are several methods to derive a POS. In the original paper by [PSCO04], two approaches are outlined to construct the ordering constraints between activities, either analyzing the resource profile to avoid all possible resource conflicts, based on Minimal Critical Sets (MCS) [LR83], or using a single-point solution together with a chaining procedure to construct resource chains (see Figure 2). They find that using the single-point solution together with chaining seems both simple and most effective. Subsequent work explored alternative MCS-based approaches [LM09, LMB13] and chaining heuristics [FLVX12].
Figure 2: Example Gannt chart (figure adjusted from [FLVX12]). The X-axis is time, and Y-axis shows resource demand. The arrows indicate an example of generating a POS with chaining, starting from a fixed solution schedule.
@@ 3.2.3 Execution Strategies
An STNU is dynamically controllable (DC) if there is a strategy to determine execution times for all controllable (non-contingent) time points which ensures that all temporal constraints are met, regardless of the outcomes of the contingent links. Some DC-checking algorithms generate new so-called wait edges to make the network DC [Mor14].
The DC STNU with wait edges is referred to as an Extended STNU (ESTNU) and is input to a Real-Time Execution Algorithm. Such an algorithm is the online component that transforms an STNU into a schedule (i.e. an execution time for each node), given the observations for the contingent nodes. An algorithm specifically tailored to ESTNUs is RTE [HP24, Pos22].
@@ 3.3 Benchmarking Approaches
Benchmarking procedures for comparing scheduling methods are inconsistent in the literature, with varying problem sets and comparison methods. Some studies used industrial scheduling instances [LM09, LMB13], while the majority [PSCO04, FLVX12, VFL16, FVL16] relied on PSPlib [KS96] instances, which are deterministic and transformed into stochastic versions with noise. Research typically focused on instances with 10, 20, and 30 activities (j10-j30). Different studies assessed varying metrics, such as schedule flexibility and robustness [PSCO04, PCOS07], solver performance [LM09, LMB13], or the $\alpha$-robust makespan [FLVX12, VFL16, FVL16]. However, no comprehensive benchmarking paper exists that evaluates both solution quality and computation time while also correctly accounting for infeasibilities. The main challenge is that not all instances can be solved by all methods, making it difficult to directly compare the distributions of solution quality or other metrics. We take inspiration from [LF03], who compare different planners that can fail, providing a framework for comparing solution quality and speed while also correctly considering these failures.
@@ 4 Scheduling Methods
This section outlines the proposed methods. We explain how to use CP for these scheduling problems so far dominated by MIP approaches. Note that in the Technical Appendix [VdHPF+24], we include a comparison between a CP and a MIP approach for RCPSP/max (CPOptimizer and CPLEX; [IBM24]) demonstrating that CP outperforms MIP, which is in line with the literature.
@@ 4.1 Constraint Programming for RCPSP/max
$\sum_{j\in J}\mathrm{Pulse}(x_{j},r_{r,j})\leq c_{r};$ $\ \forall r\in R$
$x_{j}:\mathrm{IntervalVar}(J,d_{j});$ $\ \forall j\in J$
For the deterministic RCPSP/max, we use the modern interval constraints from the IBM CP optimizer [IBM24]. In the equations above, we use IBM’s syntax and modify the RCPSP example from [LRSV18] to RCPSP/max. We use the earlier introduced nomenclature together with the minimal time lags $min_{j,i}$ and maximal time lags $max_{j,i}$ that are the temporal differences between start times of activities $j$ and $i$ if $i$ is a successor of $j$. We introduce the decision variable $x_{j}$ as the interval variable for activity $j\in J$. The $\mathrm{Pulse}$ function generates a cumulative expression over a given interval $x_{j}$ with a certain value. For a task $j$, this value is its resource usage $r_{r,j}$. The aggregated pulse values are constrained so that their sum does not exceed the total available resource capacity $c_{r}$.
@@ 4.2.1 Proactive Method
For this SAA method, we can reuse the deterministic RCPSP/max CP model, introduced in Section 4.1, but we introduce scenarios. This model is inspired by the MIP version by [VFL16]. A special variant is the SAA with only one sample, for which a $\gamma$-quantile can be used which we call $\mathsf{proactive}_{\gamma}$. If a feasible schedule can be found for the $\gamma$-quantile, this schedule will also be feasible for all duration realizations on the left-hand side of the $\gamma$-quantile because of Proposition 1. We provide the SAA model below. We use the same nomenclature as for the deterministic model, but we introduce the notion of scenarios $\omega\in\Omega$, and find a schedule $s_{j}\ \forall j\in J$ that is feasible for all scenarios it has seen if one exists:
@@
$\sum_{j\in J}\mathrm{Pulse(x_{j}^{\omega},r_{r,j})\leq c_{r}};$ $\ \forall r\in R\ \forall\omega\in\Omega$
$x_{j}^{\omega}:\mathrm{IntervalVar(J,y_{j}^{\omega})};$ $\ \forall j\in J\ \forall\omega\in\Omega$
$s_{j}=\mathrm{start}(x_{j}^{\omega});$ $\ \forall j\in J\ \forall\omega\in\Omega$
@@ 4.2.2 Reactive Method
Fully reactive approaches, involving complete rescheduling by solving a deterministic RCPSP/max, have been considered impractical due to high computational demands and low schedule stability [VdVDH07]. However, advances in CP for scheduling [LRSV18, NRR23] mitigate these issues. In the industrial setting where we are applying our work, decision-makers often reschedule their entire future plans when changes occur. Thus, including this reactive approach in our comparison is valuable. To our knowledge, such an approach has not been evaluated before. The outline is:
@@
  • At every decision moment (when an activity finishes) resolve the deterministic RCPSP/max while fixing all variables until the current time to reschedule with new information, we again use the estimation $\hat{d}$ for the activities that did not finish yet. Resolving is needed when the finish time of an activity deviates from the estimated finish time. We warm start the solver with the previous solution.
@@ 4.2.3 STNU-based Method
The resulting STNU is tested for dynamic controllability (DC), and if the network is DC its extended form (ESTNU) is given to the RTE algorithm. Since makespan minimization is of interest, we slightly adjust the algorithm from [HP24]: instead of selecting an arbitrary executable time point and an arbitrary allowed execution time, we always choose the earliest possible time point at the earliest possible execution time. In the Technical Appendix we include a step-by-step example of the STNU-based method.
@@ 4.3 Statistical Tests for Pairwise Comparison
A strategy for benchmarking is to provide partial orderings of the scheduling methods for the different metrics of interest (e.g. solution quality, runtime offline, runtime online). A partial ordering can be obtained by executing pairwise comparisons of the methods per problem size and per metric, taking inspiration from [LF03].
@@
An alternative test to the Wilcoxon test that is also used by [LF03] is the proportion test (see Test 4 in the book by [Kan06]). This test is weaker than the Wilcoxon, but provides at least information about significance in the proportion of wins when the Wilcoxon test shows no significant difference. The magnitude test provides more insight into the magnitude differences of the performance metrics. This test is also known as the pairwise t-test on two related samples of scores (see Test 10 in the book by [Kan06]). This test can only be performed on double hits because infinitely bad computation time or solution quality will disturb the test.
@@
Table 1: Pairwise test results for ubo50, $\epsilon=1$, covering solution quality and runtime. Metrics include Wilcoxon, proportion, and magnitude tests (Section 4.3). Results: [pairs] z-value (p-value) (* for $p<0.05$), proportion (p-value), and t-stat (p-value) with normalized averages. Additional results for other $\epsilon$, j10–30, ubo50, and ub0100 are in the Technical Appendix. Column headers list the compared methods, excluding zero-difference pairs.
@@ 5 Experimental Evaluation
The goal of the evaluation is to analyze the relative performance of the different proposed scheduling methods.
@@ 5.1 Data Generation
We use the j10, j20, j30, ubo50, and ubo100 sets from the PSPlib [KS96]. Instead of using all instances from j10 to j30, we select 50 per set and extend previous work [VFL16, FVL16] by including 50 instances each from the ubo50 and ubo100 sets. Following prior research, we use deterministic durations to set up stochastic distributions with different noise levels, converting deterministic instances into stochastic ones. We use uniform discrete distributions for durations based on the deterministic processing times $d_{j}$. Specifically, we define the lower bound as $lb_{j}=\max(1,\text{round}(d_{j}-\epsilon\cdot\sqrt{d_{j}}))$ and the upper bound as $ub_{j}=\text{round}(d_{j}+\epsilon\cdot\sqrt{d_{j}})$, where we vary the noise level $\epsilon=\{1,2\}$. The source and sink nodes always have a deterministic duration of zero. Each evaluation of a method corresponds to one sample from the distribution, and we sample $10$ times for each instance. We excluded the instance samples for which it is impossible to find a feasible schedule (i.e. the deterministic problem with perfect information is infeasible for the given activity duration sample).